# MPG-SAM 2: Adapting SAM 2 with Mask Priors and Global Context for Referring Video Object Segmentation

### v3

@@ 1 Introduction
Referring video object segmentation (RVOS) [45, 30, 27, 21] aims to segment the target objects in a video according to textual descriptions. This task, combining video segmentation and language comprehension, necessitates not only proficiency in segmentation and inter-frame information propagation, as in traditional video object segmentation (VOS) [32, 6, 1], but also a robust understanding of the textual reference within the broader video context. The core challenges of RVOS thus lie in efficiently aligning multimodal information and maintaining temporal consistency.
@@ 2 Related work
Referring Video Object Segmentation. RVOS attracts extensive interest as it bridges visual and linguistic domains. Early methods, such as RefVOS [2], consider RVOS as an expansion of referring image segmentation (RIS) to the video domain. URVOS [37] advances this by combining RIS and semi-supervised video object segmentation within a unified framework using attention mechanisms. In subsequent studies [29, 40, 22, 12, 43], researchers emphasize cross-modal interaction [13, 46], further enhancing RVOS performance. However, despite notable gains, the computational expense and complexity of multi-stage pipelines limit practical feasibility. In response, query-based Transformer architectures offer efficient solutions with simplified yet robust frameworks. Notably, MTTR [3] and ReferFormer [45] pioneer the use of the DETR series [5, 57] in RVOS, introducing novel multimodal interaction mechanisms. Recent methods [27, 30, 38, 21, 33] refine these Transformer architectures through advanced temporal and multimodal feature integration techniques. For instance, SgMg [30] enhances ReferFormer by replacing dynamic convolution with a segmentation optimizer and uses spectral information to guide visual feature fusion. SOC [27] implements layered temporal modeling across video and object levels, achieving early multimodal fusion. More recently, Losh [53] introduces a joint prediction network for short and long sentences, reducing the over-influence of action and relational cues in segmentation. VD-IT [58] explores the potential of text-to-video diffusion models in the RVOS task, leveraging the inherent rich semantics and coherent temporal correspondences of video generation models to ensure temporal instance consistency. To tackle longer sequences and complex scenarios in the MeViS [11] dataset, DsHmp [15] proposes a static text decoupling strategy, enhancing the temporal understanding of static and dynamic content at both frame and object levels, thereby capturing both short-term and long-term motion insights.
@@
Segment Anything Model. SAM [19] is an interactive segmentation model capable of generating non-semantic masks based on various prompts. Trained on a large-scale dataset, SAM exhibits strong generalization across a wide range of common objects. Several variants [47, 55, 17] have focused on improving the segmentation accuracy and computational efficiency of SAM. Moreover, SAM has found widespread application in various fields, including video tracking [8], remote sensing image interpretation [39], and medical image processing [54]. Recently, SAM 2 [17] extends SAM to the video domain, achieving state-of-the-art performance. Despite its excellence in visual segmentation tasks using box, point, or mask prompts, SAM lacks language comprehension capabilities and cannot directly handle referring segmentation tasks. With the advancement of multimodal large language models (MLLMs), recent approaches [20, 48, 35] utilize MLLMs to encode textual guidance for SAM segmentation. Specifically, LISA [20] fine-tunes LLaVA [24] to generate multimodal features by extracting hidden embeddings based on specific prompts. u-LLaVA [48] extends this approach to enable joint multi-task processing at region and pixel levels. GLaMM [35] integrates the visual grounding task and incorporates language responses to provide multi-granular segmentation prompts to the model. Most recently, EVF-SAM [56] introduces a lightweight pre-fusion architecture that employs joint visual-language encoding to generate high-quality textual prompts, leading to excellent segmentation performance.
@@ 4.4 Model Analysis
Additionally, we explore different configurations for the global video feature input to the hierarchical global-historical aggregator. One straightforward approach named “Vanilla” employs the video embeddings directly output by the multimodal encoder as the global feature. An alternative approach dubbed “Masked” leverages a video feature enriched with mask prior information. Results in Tab. 4 indicate that video features containing mask information are more beneficial for global context integration, as they better emphasize global frame mask information to guide the segmentation of the current frame.

### v4

@@ 1 Introduction
Referring video object segmentation (RVOS) [48, 32, 29, 23] aims to segment the target objects in a video according to textual descriptions. This task, combining video segmentation [21] and language comprehension, necessitates not only proficiency in segmentation and inter-frame information propagation, as in traditional video object segmentation (VOS) [34, 6, 1], but also a robust understanding of the textual reference within the broader video context. The core challenges of RVOS thus lie in efficiently aligning multimodal information and maintaining temporal consistency.
@@ 2 Related work
Referring Video Object Segmentation. RVOS attracts extensive interest as it bridges visual and linguistic domains. Early methods, such as RefVOS [2], consider RVOS as an expansion of referring image segmentation (RIS) to the video domain. URVOS [40] advances this by combining RIS and semi-supervised video object segmentation within a unified framework using attention mechanisms. In subsequent studies [31, 43, 24, 12, 46], researchers emphasize cross-modal interaction [13, 49], further enhancing RVOS performance. However, despite notable gains, the computational expense and complexity of multi-stage pipelines limit practical feasibility. In response, query-based Transformer architectures [22] offer efficient solutions with simplified yet robust frameworks. Notably, MTTR [3] and ReferFormer [48] pioneer the use of the DETR series [5, 60] in RVOS, introducing novel multimodal interaction mechanisms. Recent methods [29, 32, 41, 23, 35] refine these Transformer architectures through advanced temporal and multimodal feature integration techniques. For instance, SgMg [32] enhances ReferFormer by replacing dynamic convolution with a segmentation optimizer and uses spectral information to guide visual feature fusion. SOC [29] implements layered temporal modeling across video and object levels, achieving early multimodal fusion. More recently, Losh [56] introduces a joint prediction network for short and long sentences, reducing the over-influence of action and relational cues in segmentation. VD-IT [61] explores the potential of text-to-video diffusion models in the RVOS task, leveraging the inherent rich semantics and coherent temporal correspondences of video generation models to ensure temporal instance consistency. To tackle longer sequences and complex scenarios in the MeViS [11] dataset, DsHmp [15] proposes a static text decoupling strategy, enhancing the temporal understanding of static and dynamic content at both frame and object levels, thereby capturing both short-term and long-term motion insights.
@@
Segment Anything Model. SAM [19] is an interactive segmentation model capable of generating non-semantic masks based on various prompts. Trained on a large-scale dataset, SAM exhibits strong generalization across a wide range of common objects. Several variants [50, 58, 17] have focused on improving the segmentation accuracy and computational efficiency of SAM. Moreover, SAM has found widespread application in various fields, including video tracking [8], remote sensing image interpretation [42, 39], and medical image processing [57]. Recently, SAM 2 [17] extends SAM to the video domain, achieving state-of-the-art performance. Despite its excellence in visual segmentation tasks using box, point, or mask prompts, SAM lacks language comprehension capabilities and cannot directly handle referring segmentation tasks. With the advancement of multimodal large language models (MLLMs), recent approaches [20, 51, 37] utilize MLLMs to encode textual guidance for SAM segmentation. Specifically, LISA [20] fine-tunes LLaVA [26] to generate multimodal features by extracting hidden embeddings based on specific prompts. u-LLaVA [51] extends this approach to enable joint multi-task processing at region and pixel levels. GLaMM [37] integrates the visual grounding task and incorporates language responses to provide multi-granular segmentation prompts to the model. Most recently, EVF-SAM [59] introduces a lightweight pre-fusion architecture that employs joint visual-language encoding to generate high-quality textual prompts, leading to excellent segmentation performance.
@@ 4.4 Model Analysis
Additionally, we explore different configurations for the global video feature input to the hierarchical global-historical aggregator. One straightforward approach named “Vanilla" employs the video embeddings directly output by the multimodal encoder as the global feature. An alternative approach dubbed “Masked" leverages a video feature enriched with mask prior information. Results in Tab. 4 indicate that video features containing mask information are more beneficial for global context integration, as they better emphasize global frame mask information to guide the segmentation of the current frame.
@@ Appendix B Additional Visualization Results
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62431020, in part by the Foundation for Innovative Research Groups of Hubei Province under Grant 2024AFA017, in part by the Fundamental Research Funds for the Central Universities under Grant 2042025kf0030.